an exercise machine with an original price of $950 is on sale at 20% off.
what is the discount?
what is the machine's sale price?

Answer by TimothyLamb(4379)

discount = 950*0.20 = $190
---
sale price = 950*(1 - 0.20) = $760
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---